# Robust Point Matching with Distance Profiles

### v2

@@ 2 Outlier Robustness of Distance Profile Matching
In other words, Assumption 1 is equivalent to letting $\mu=\lambda\rho+(1-\lambda)\mu_{o}$ and $\nu=\lambda\rho+(1-\lambda)\nu_{o}$, where the common part $\lambda\rho$ is the mixture consisting of the first $K$ components and the outlier parts $(1-\lambda)\mu_{o},(1-\lambda)\nu_{o}$ are the mixtures of the remaining components. The motivation behind Assumption 1 is that the objects $\mu,\nu$ are composed of several regions as we have seen from the example in Section 1.1, where each region is represented as a Gaussian distribution so that the mean corresponds to the center of the region and the variance roughly represents the size of the region. The points corresponding to the same region are to be matched, while we ignore matching the points from the outlier components.
@@
There are several factors determining the success of the matching. First, the objects $\mu,\nu$ must be well-configured in that the centers avoid certain symmetries that cannot be distinguished by any procedures based on pairwise distances. For instance, suppose $K=3$ and $\|\theta_{k}-\theta_{k^{\prime}}\|_{2}$ is the same for any $k\neq k^{\prime}$, meaning that all components are equidistant from each other. Then, any matching procedure based on $\|X_{i}-X_{\ell}\|_{2}$’s and $\|Y_{j}-Y_{\ell}\|_{2}$’s, such as the proposed one or the usual quadratic matching, essentially treats all the points as the same point and thus fail to distinguish the components. Therefore, we need a suitable assumption on the configuration of the centers for matching to be well-defined. In addition to proper configurations, the sample sizes $n,m$ must be sufficiently large to approximate $\mu,\nu$ by the samples. Meanwhile, the variances $\sigma_{k}$’s and $\tau_{k}$’s must be sufficiently small to ensure that the mixture components are well-separated from each other. The sample sizes also affect how small the variances should be. In fact, the larger the sample sizes $n,m$ are, the smaller the variances should be to uniformly control the deviation of the samples from their centers. The following theorem formally shows the matching guarantee where the aforementioned conditions on the configurations, sample sizes, and variances are stated as (i), (ii), and (iii), respectively.
@@
Suppose $\mu$ and $\nu$ satisfy Assumption 1. Let $X_{1},\ldots,X_{n}$ and $Y_{1},\ldots,Y_{m}$ be independent samples from $\mu$ and $\nu$, respectively. Let $\pi\colon[n]\to[m]$ be the output of Algorithm 1 with $\mathcal{D}=W_{1}$. For $i\in[n]$ and $j\in[m]$, let $t(i)\in[t]$ and $s(j)\in[s]$ denote the components where $X_{i}$ and $Y_{j}$ are from, respectively. Then, for $\delta\in(0,1)$ such that $\delta\leq 12e^{-4(t\vee s)/5}$, we have
@@
if the following conditions hold.
  • (i)
    Letting
@@
    $\overline{W}(\alpha,\beta):=W_{1}\left(\sum_{k=1}^{t}p_{k}\delta_{\|\theta_{\alpha}-\theta_{k}\|_{2}},\sum_{k=1}^{s}q_{k}\delta_{\|\eta_{\beta}-\eta_{k}\|_{2}}\right)\quad\forall(\alpha,\beta)\in[t]\times[s],$
    suppose
    $\min_{\alpha\neq\beta}\overline{W}(\alpha,\beta)-\max_{\alpha\in[K]}\overline{W}(\alpha,\alpha)=:\Omega>0.$ (2.2)
  • (ii)
    Letting $R=\max_{\alpha,\beta\in[t]}\|\theta_{\alpha}-\theta_{\beta}\|_{2}\vee\max_{\alpha,\beta\in[s]}\|\eta_{\alpha}-\eta_{\beta}\|_{2}$, suppose
    $n\wedge m\geq\frac{1600R^{2}\log(12/\delta)}{\Omega^{2}}.$ (2.3)
  • (iii)
    Letting $\sigma=\max_{k\in[t]}\sigma_{k}$ and $\tau=\max_{k\in[s]}\tau_{k}$, suppose
    $\sigma^{2}\leq\frac{\Omega^{2}}{256(d\vee 8\log(4n^{2}/\delta))}\quad\text{and}\quad\tau^{2}\leq\frac{\Omega^{2}}{256(d\vee 8\log(4m^{2}/\delta))}.$ (2.4)
As discussed earlier, the condition (i) of Theorem 1 ensures that the centers of the mixture components are well-configured so that the matching is well-defined. Particularly, (2.2) ensures that for any $\alpha\in[K]$,
$\operatorname*{arg\,min}_{\beta\in[s]}\overline{W}(\alpha,\beta)=\alpha,$
implying that $\eta_{\alpha}$ is closest to $\theta_{\alpha}$ in terms of the distance profiles $\{\|\theta_{\alpha}-\theta_{k}\|_{2}\}_{k=1}^{t}$ and $\{\|\eta_{\alpha}-\eta_{k}\|_{2}\}_{k=1}^{s}$ properly weighted by $(p_{1},\ldots,p_{t})$ and $(q_{1},\ldots,q_{s})$, respectively, under $W_{1}$. The condition (ii) ensures that the sample sizes $n,m$ are sufficiently large to approximate $\mu,\nu$. The condition (iii) characterizes how small the variances $\sigma_{k}$’s and $\tau_{k}$’s should be given the sample sizes $n,m$ and $\Omega$.
Proof of Theorem 1.
Note that $D(i,j)$ of Algorithm 1 is $W(i,j)$ defined in (1.1) as we let $\mathcal{D}=W_{1}$. For any pair $(i,j)\in[n]\times[m]$, we show that $W(i,j)\approx\overline{W}(t(i),s(j))$ holds. To this end, we define an intermediate quantity $W^{\prime}(i,j)$ as follows: for any $(i,j)\in[n]\times[m]$,
@@
$\begin{split}W^{\prime}(i,j)&:=W_{1}\left(\frac{1}{n}\sum_{\ell=1}^{n}\delta_{\|\theta_{t(i)}-\theta_{t(\ell)}\|_{2}},\frac{1}{m}\sum_{\ell=1}^{m}\delta_{\|\eta_{s(j)}-\eta_{s(\ell)}\|_{2}}\right)\\ &=W_{1}\left(\sum_{k=1}^{t}\frac{n_{k}}{n}\delta_{\|\theta_{t(i)}-\theta_{k}\|_{2}},\sum_{k=1}^{s}\frac{m_{k}}{m}\delta_{\|\eta_{s(j)}-\eta_{k}\|_{2}}\right),\end{split}$
where we define
$n_{k}=\sum_{\ell=1}^{n}1\{t(\ell)=k\}\quad\text{and}\quad m_{k}=\sum_{\ell=1}^{m}1\{s(\ell)=k\}.$
We first show that $W^{\prime}(i,j)\approx\overline{W}(t(i),s(j))$ for any $(i,j)\in[n]\times[m]$. We use the fact—see, for instance, Section 2.2 of [Villani, 2003]—that the Wasserstein-1 distance on $\mathscr{P}(\mathbb{R})$ can be written as the $L^{1}$ distance between cumulative distribution functions, which yields
$\overline{W}(\alpha,\beta)$ $=\int_{\mathbb{R}}\left|\sum_{k=1}^{t}p_{k}1\{\|\theta_{\alpha}-\theta_{k}\|_{2}\leq z\}-\sum_{k=1}^{s}q_{k}1\{\|\eta_{\beta}-\eta_{k}\|_{2}\leq z\}\right|\,\mathrm{d}z\quad\forall(\alpha,\beta)\in[t]\times[s],$
$W^{\prime}(i,j)$ $=\int_{\mathbb{R}}\left|\sum_{k=1}^{t}\frac{n_{k}}{n}1\{\|\theta_{t(i)}-\theta_{k}\|_{2}\leq z\}-\sum_{k=1}^{s}\frac{m_{k}}{m}1\{\|\eta_{s(j)}-\eta_{k}\|_{2}\leq z\}\right|\,\mathrm{d}z\quad\forall(i,j)\in[n]\times[m].$
Hence,
$\begin{split}&|W^{\prime}(i,j)-\overline{W}(t(i),s(j))|\\ &\leq\int_{\mathbb{R}}\left|\sum_{k=1}^{t}\left(\frac{n_{k}}{n}-p_{k}\right)1\{\|\theta_{t(i)}-\theta_{k}\|_{2}\leq z\}\right|\,\mathrm{d}z+\int_{\mathbb{R}}\left|\sum_{k=1}^{s}\left(\frac{m_{k}}{m}-q_{k}\right)1\{\|\eta_{s(j)}-\eta_{k}\|_{2}\leq z\}\right|\,\mathrm{d}z\\ &=\int_{0}^{R}\left|\sum_{k=1}^{t}\left(\frac{n_{k}}{n}-p_{k}\right)1\{\|\theta_{t(i)}-\theta_{k}\|_{2}\leq z\}\right|\,\mathrm{d}z+\int_{0}^{R}\left|\sum_{k=1}^{s}\left(\frac{m_{k}}{m}-q_{k}\right)1\{\|\eta_{s(j)}-\eta_{k}\|_{2}\leq z\}\right|\,\mathrm{d}z\\ &\leq R\left(\sum_{k=1}^{t}\left|\frac{n_{k}}{n}-p_{k}\right|+\sum_{k=1}^{s}\left|\frac{m_{k}}{m}-q_{k}\right|\right)=:R(Z_{1}+Z_{2}).\end{split}$
Next, we show $W(i,j)\approx W^{\prime}(i,j)$ for any $(i,j)\in[n]\times[m]$. By the triangle inequality of $W_{1}$, we have
@@
$\begin{split}&|W(i,j)-W^{\prime}(i,j)|\\ &\leq W_{1}\left(\frac{1}{n}\sum_{\ell=1}^{n}\delta_{\|X_{i}-X_{\ell}\|_{2}},\frac{1}{n}\sum_{\ell=1}^{n}\delta_{\|\theta_{t(i)}-\theta_{t(\ell)}\|_{2}}\right)+W_{1}\left(\frac{1}{m}\sum_{\ell=1}^{m}\delta_{\|Y_{j}-Y_{\ell}\|_{2}},\frac{1}{m}\sum_{\ell=1}^{m}\delta_{\|\eta_{s(j)}-\eta_{s(\ell)}\|_{2}}\right).\end{split}$
Note that we can write $X_{i}=\theta_{t(i)}+\xi_{i}$ and $Y_{j}=\eta_{s(j)}+\zeta_{j}$ for any $(i,j)\in[n]\times[m]$, where $\xi_{i}\,|\,t(i)\sim N(0,\sigma_{t(i)}^{2}I_{d})$ and $\zeta_{j}\,|\,s(j)\sim N(0,\tau_{s(j)}^{2}I_{d})$ are independent. Recall that the Wasserstein-1 distance between two sets of points in $\mathbb{R}$ with the same cardinality satisfies the following:
@@
$W_{1}\left(\frac{1}{n}\sum_{\ell=1}^{n}\delta_{\|X_{i}-X_{\ell}\|_{2}},\frac{1}{n}\sum_{\ell=1}^{n}\delta_{\|\theta_{t(i)}-\theta_{t(\ell)}\|_{2}}\right)\leq\frac{1}{n}\sum_{\ell=1}^{n}\left|\|X_{i}-X_{\ell}\|_{2}-\|\theta_{t(i)}-\theta_{t(\ell)}\|_{2}\right|.$
Since $|\|X_{i}-X_{\ell}\|_{2}-\|\theta_{t(i)}-\theta_{t(\ell)}\|_{2}=|\|\theta_{t(i)}-\theta_{t(\ell)}+\xi_{i}-\xi_{\ell}\|_{2}-\|\theta_{t(i)}-\theta_{t(\ell)}\|_{2}|\leq\|\xi_{i}-\xi_{\ell}\|_{2}$, we have
$W_{1}\left(\frac{1}{n}\sum_{\ell=1}^{n}\delta_{\|X_{i}-X_{\ell}\|_{2}},\frac{1}{n}\sum_{\ell=1}^{n}\delta_{\|\theta_{t(i)}-\theta_{t(\ell)}\|_{2}}\right)\leq\frac{1}{n}\sum_{\ell=1}^{n}\|\xi_{i}-\xi_{\ell}\|_{2}\leq\max_{i,\ell\in[n]}\|\xi_{i}-\xi_{\ell}\|_{2}.$

### v5

@@ 1.2 Stylized Example
Figure 1(c) and Figure 1(f) repeat the same procedure after removing 16 points around the fin of the target dolphin in Figure 1(e) and clustering again with $k=6$. Figure 1(c) shows that the matching result is similar to the previous one. In this case, though the points around the fin of the source dolphin do not have corresponding points in the target dolphin and thus we may ignore matching them, the proposed procedure still matches them to the points around the peduncle (magenta) of the target dolphin, which is a cluster adjacent to the removed fin. This stylized example confirms the performance of the proposed distance profile matching procedure in practice.
@@ 2 Outlier Robustness of Distance Profile Matching
Assumption 1 is equivalent to letting $\mu=\lambda\rho+(1-\lambda)\mu_{o}$ and $\nu=\lambda\rho+(1-\lambda)\nu_{o}$, where the common part $\lambda\rho$ is the mixture consisting of the first $K$ components and the other parts, namely, the outlier parts $(1-\lambda)\mu_{o},(1-\lambda)\nu_{o}$ are the mixtures of the remaining components. The motivation behind Assumption 1 is that the objects $\mu,\nu$ are composed of several regions as we have seen from the example in Section 1.1, where each region is represented as a Gaussian distribution so that the mean corresponds to the center of the region and the variance roughly represents the size of the region. The points corresponding to the same region are to be matched, while we ignore matching the points from the outlier components.
@@
Suppose $\mu$ and $\nu$ satisfy Assumption 1. Let $X_{1},\ldots,X_{n}$ and $Y_{1},\ldots,Y_{m}$ be independent samples from $\mu$ and $\nu$, respectively. Let $\pi\colon[n]\to[m]$ be the output of Algorithm 1 with $\mathcal{D}=W_{1}$. For $i\in[n]$ and $j\in[m]$, let $t(i)\in[t]$ and $s(j)\in[s]$ denote the components where $X_{i}$ and $Y_{j}$ are from, respectively. Then, for any $\delta\in(0,1)$, we have
@@
if
$\begin{split}&\min_{\begin{subarray}{c}\alpha\in[K]\\ \beta\in[s]\backslash\{\alpha\}\end{subarray}}\overline{W}(\alpha,\beta)-R\cdot\bigg{(}1-\sum_{k\in[K]}p_{k}\bigg{)}\\ &\geq\left(16\Gamma\sqrt{d}\right)\vee\bigg{(}8R\sqrt{\frac{\log K+\log(t+s-2)+\log\tfrac{4}{\delta}}{2(n\wedge m)}}\bigg{)}\vee\bigg{(}32\Gamma\sqrt{2\log\frac{4(n\vee m)^{2}}{\delta}}\bigg{)},\end{split}$ (2.3)
where $R=\max\{\|x-y\|_{2}:x,y\in\{\theta_{1},\ldots,\theta_{t},\eta_{1},\ldots,\eta_{s}\}\}$, $\Gamma=\max_{k\in[t]}\sigma_{k}\vee\max_{k\in[s]}\tau_{k}$, and
@@
$\overline{W}(\alpha,\beta):=W_{1}\left(\sum_{k=1}^{t}p_{k}\delta_{\|\theta_{\alpha}-\theta_{k}\|_{2}},\sum_{k=1}^{s}q_{k}\delta_{\|\eta_{\beta}-\eta_{k}\|_{2}}\right)\quad\forall(\alpha,\beta)\in[t]\times[s].$ (2.4)
To understand Theorem 1, it is instructive to begin with the clean case where the noise is absent, that is, $\Gamma=0$, and there is no sampling error, namely, exactly $np_{k}$ and $mq_{k}$ points among $X_{i}$’s and $Y_{j}$’s are from the $k$-th component of $\mu$ and $\nu$, respectively. Then, the point clouds are essentially the sets of centers $\{\theta_{1},\ldots,\theta_{t}\}$ and $\{\eta_{1},\ldots,\eta_{s}\}$ weighted by $(p_{1},\ldots,p_{t})$ and $(q_{1},\ldots,q_{s})$, respectively. Therefore, the Wasserstein-1 distances between distance profiles (1.1) boil down to the Wasserstein-1 distance between the weighted sets of centers (2.4). Then, the success of Algorithm 1, namely, the event (2.1), is equivalent to having $\operatorname*{arg\,min}_{\beta\in[s]}\overline{W}(\alpha,\beta)=\alpha$ for any $\alpha\in[K]$, which is guaranteed by having
$\min_{\beta\in[s]\backslash\{\alpha\}}\overline{W}(\alpha,\beta)>\overline{W}(\alpha,\alpha).$ (2.5)
To ensure this for all $\alpha\in[K]$ simultaneously, we require a stronger condition, [^2] Note that (2.6) is much stronger than requiring (2.5) for all $\alpha\in[K]$. We chose to present Theorem 1 based on (2.6) as it leads to a more interpretable condition in terms of the outlier proportion. namely,
$\min_{\begin{subarray}{c}\alpha\in[K]\\ \beta\in[s]\backslash\{\alpha\}\end{subarray}}\overline{W}(\alpha,\beta)-\max_{\alpha\in[K]}\overline{W}(\alpha,\alpha)>0.$ (2.6)
How can this condition be satisfied? First, suppose there is no outlier at all, namely, $K=t=s$. Then, since $\overline{W}(\alpha,\alpha)=0$ for any $\alpha\in[K]$, it suffices to have $\overline{W}(\alpha,\beta)>0$ for any $\beta\in[s]\backslash\{\alpha\}$, which can be guaranteed if the pairwise distances among centers are well-separated. In general, we can upper bound $\overline{W}(\alpha,\alpha)$ for any $\alpha\in[K]$, which is essentially controlled by the outlier proportion $1-\sum_{k\in[K]}p_{k}$, as we show in Lemma 1. Meanwhile, for $\alpha\in[K]$ and $\beta\in[s]\backslash\{\alpha\}$, the quantity $\overline{W}(\alpha,\beta)$ is essentially of a constant order, namely, $O(1)$. Therefore, the condition (2.6) is satisfied if the outlier proportion is small enough, which is captured by the left-hand side of (2.3).
In general cases with $\Gamma>0$, where the point clouds are samples, the Wasserstein-1 distances (1.1) deviate from (2.4) due to the noise and sampling error. Therefore, the condition (2.6) needs to be modified to reflect this deviation, and characterizing this probabilistic deviation leads to the right-hand side of (2.3).
Based on the above insights, we prove Theorem 1 as follows.
Proof.
We analyze the deviation of (1.1) from (2.4), namely, for any pair $(i,j)\in[n]\times[m]$, we derive an upper bound on the deviation $|W(i,j)-\overline{W}(t(i),s(j))|$. To this end, we define an intermediate quantity $W^{\prime}(i,j)$ as follows: for any $(i,j)\in[n]\times[m]$,
$\begin{split}W^{\prime}(i,j)&:=W_{1}\bigg{(}\frac{1}{n}\sum_{\ell=1}^{n}\delta_{\|\theta_{t(i)}-\theta_{t(\ell)}\|_{2}},\frac{1}{m}\sum_{\ell=1}^{m}\delta_{\|\eta_{s(j)}-\eta_{s(\ell)}\|_{2}}\bigg{)}\\ &=W_{1}\bigg{(}\sum_{k=1}^{t}\frac{n_{k}}{n}\delta_{\|\theta_{t(i)}-\theta_{k}\|_{2}},\sum_{k=1}^{s}\frac{m_{k}}{m}\delta_{\|\eta_{s(j)}-\eta_{k}\|_{2}}\bigg{)},\end{split}$
where we define $n_{k}=\sum_{\ell=1}^{n}1\{t(\ell)=k\}$ and $m_{k}=\sum_{\ell=1}^{m}1\{s(\ell)=k\}$. We first derive an upper bound on $|W^{\prime}(i,j)-\overline{W}(t(i),s(j))|$ for any $(i,j)\in[n]\times[m]$. By the triangle inequality of $W_{1}$, we have
@@
$\begin{split}&|W^{\prime}(i,j)-\overline{W}(t(i),s(j))|\\ &\leq\underbrace{W_{1}\bigg{(}\sum_{k=1}^{t}\frac{n_{k}}{n}\delta_{\|\theta_{t(i)}-\theta_{k}\|_{2}},\sum_{k=1}^{t}p_{k}\delta_{\|\theta_{t(i)}-\theta_{k}\|_{2}}\bigg{)}}_{=:(*)}+\underbrace{W_{1}\bigg{(}\sum_{k=1}^{s}\frac{m_{k}}{m}\delta_{\|\eta_{s(j)}-\eta_{k}\|_{2}},\sum_{k=1}^{s}q_{k}\delta_{\|\eta_{s(j)}-\eta_{k}\|_{2}}\bigg{)}}_{=:(**)}.\end{split}$
For each $\alpha\in[K]$, order $\theta_{1},\ldots,\theta_{t}$ by the distance to $\theta_{\alpha}$, namely, find $k_{\alpha}\colon[t]\to[t]$ such that $\|\theta_{\alpha}-\theta_{k_{\alpha}(1)}\|_{2}\leq\cdots\leq\|\theta_{\alpha}-\theta_{k_{\alpha}(t)}\|_{2}$. Then, by Lemma 2, we have
@@
$\begin{split}(*)&\leq\max_{\alpha\in[K]}W_{1}\bigg{(}\sum_{k=1}^{t}\frac{n_{k}}{n}\delta_{\|\theta_{\alpha}-\theta_{k}\|_{2}},\sum_{k=1}^{t}p_{k}\delta_{\|\theta_{\alpha}-\theta_{k}\|_{2}}\bigg{)}\\ &\leq\max_{\alpha\in[K]}\|\theta_{\alpha}-\theta_{k_{\alpha}(t)}\|_{2}\cdot\max_{j\in[t-1]}\left|\sum_{k=1}^{j}\left(\frac{n_{k_{\alpha}(k)}}{n}-p_{k_{\alpha}(k)}\right)\right|\\ &\leq R\cdot\max_{\alpha\in[K]}\max_{j\in[t-1]}\left|\sum_{k=1}^{j}\left(\frac{n_{k_{\alpha}(k)}}{n}-p_{k_{\alpha}(k)}\right)\right|=:R\cdot Z_{1}.\end{split}$
Similarly, for each $\beta\in[K]$, order $\eta_{1},\ldots,\eta_{s}$ by the distance to $\eta_{\beta}$, namely, find $\ell_{\beta}\colon[s]\to[s]$ such that $\|\eta_{\beta}-\eta_{\ell_{\beta}(1)}\|_{2}\leq\cdots\leq\|\eta_{\beta}-\eta_{\ell_{\beta}(s)}\|_{2}$. Then, by Lemma 2, we have
@@
$\begin{split}(**)&\leq\max_{\beta\in[K]}W_{1}\bigg{(}\sum_{k=1}^{s}\frac{m_{k}}{m}\delta_{\|\eta_{\beta}-\eta_{k}\|_{2}},\sum_{k=1}^{s}q_{k}\delta_{\|\eta_{\beta}-\eta_{k}\|_{2}}\bigg{)}\\ &\leq\max_{\beta\in[K]}\|\eta_{\beta}-\eta_{\ell_{\beta}(s)}\|_{2}\cdot\max_{j\in[s-1]}\left|\sum_{k=1}^{j}\left(\frac{m_{\ell_{\beta}(k)}}{m}-q_{\ell_{\beta}(k)}\right)\right|\\ &\leq R\cdot\max_{\beta\in[K]}\max_{j\in[s-1]}\left|\sum_{k=1}^{j}\left(\frac{m_{\ell_{\beta}(k)}}{m}-q_{\ell_{\beta}(k)}\right)\right|=:R\cdot Z_{2}.\end{split}$
Hence, $|W^{\prime}(i,j)-\overline{W}(t(i),s(j))|\leq R\cdot(Z_{1}+Z_{2})$. Next, we bound $|W(i,j)-W^{\prime}(i,j)|$ for any $(i,j)\in[n]\times[m]$. By the triangle inequality again, we have
$\begin{split}&|W(i,j)-W^{\prime}(i,j)|\\ &\leq W_{1}\bigg{(}\frac{1}{n}\sum_{\ell=1}^{n}\delta_{\|X_{i}-X_{\ell}\|_{2}},\frac{1}{n}\sum_{\ell=1}^{n}\delta_{\|\theta_{t(i)}-\theta_{t(\ell)}\|_{2}}\bigg{)}+W_{1}\bigg{(}\frac{1}{m}\sum_{\ell=1}^{m}\delta_{\|Y_{j}-Y_{\ell}\|_{2}},\frac{1}{m}\sum_{\ell=1}^{m}\delta_{\|\eta_{s(j)}-\eta_{s(\ell)}\|_{2}}\bigg{)}.\end{split}$
Meanwhile, notice that $X_{i}=\theta_{t(i)}+\xi_{i}$ and $Y_{j}=\eta_{s(j)}+\zeta_{j}$ for any $(i,j)\in[n]\times[m]$, where $\xi_{i}\,|\,t(i)\sim N(0,\sigma_{t(i)}^{2}I_{d})$ and $\zeta_{j}\,|\,s(j)\sim N(0,\tau_{s(j)}^{2}I_{d})$ are independent. Recall that the Wasserstein-1 distance between two sets of points in $\mathbb{R}$ with the same cardinality satisfies the following:
$W_{1}\bigg{(}\frac{1}{n}\sum_{\ell=1}^{n}\delta_{\|X_{i}-X_{\ell}\|_{2}},\frac{1}{n}\sum_{\ell=1}^{n}\delta_{\|\theta_{t(i)}-\theta_{t(\ell)}\|_{2}}\bigg{)}\leq\frac{1}{n}\sum_{\ell=1}^{n}\left|\|X_{i}-X_{\ell}\|_{2}-\|\theta_{t(i)}-\theta_{t(\ell)}\|_{2}\right|.$
From $|\|X_{i}-X_{\ell}\|_{2}-\|\theta_{t(i)}-\theta_{t(\ell)}\|_{2}=|\|\theta_{t(i)}-\theta_{t(\ell)}+\xi_{i}-\xi_{\ell}\|_{2}-\|\theta_{t(i)}-\theta_{t(\ell)}\|_{2}|\leq\|\xi_{i}-\xi_{\ell}\|_{2}$, we have
$W_{1}\bigg{(}\frac{1}{n}\sum_{\ell=1}^{n}\delta_{\|X_{i}-X_{\ell}\|_{2}},\frac{1}{n}\sum_{\ell=1}^{n}\delta_{\|\theta_{t(i)}-\theta_{t(\ell)}\|_{2}}\bigg{)}\leq\frac{1}{n}\sum_{\ell=1}^{n}\|\xi_{i}-\xi_{\ell}\|_{2}\leq\max_{i,\ell\in[n]}\|\xi_{i}-\xi_{\ell}\|_{2}=:U.$